2. Suppose that you are given the following data for the country of Donaldland:

            consumption             14,000               net factor income from abroad              300

            wages                        11,000               interest payments                                2,000

            investment                  3,600               capital consumption allowance            3,300

            exports                        2,500               government spending on goods            3,600

            imports                       3,100                   and services

            profits                         3,700               rental payments                                     800

From the data provided, find

            a. GDP using the expenditure approach.

            b. GNP and NNP.

            c. NI using the income approach.

            d. statistical discrepancy.

            e. GDP using the income approach.
Answer #1

Ans a ) expenditure method : private consumption expenditure+ govt final consumption expenditure+ gross fixed capital formation + change in stocks + Net exports.

= 14000+ 3600+ 3600+ 2500- 3100

= 20600

GDP = 20600

To get GNP we deduct capital consumption allowance from this:

So

GNP= 20600-3300

= 17300

NNP = GNP + net capital from abroad

= 17300+ 300

= 17600

Income method : wages + interest payment+ profit + rental payment

= 11000+2000+3700+800

= 17500

NDP = 17500

NI by income method = NDP + net income from abroad

= 17500+ 300

= 17800

GDP by income method : NDP + capital consumption net income from abroad

= 17500+ 3300- 300

= 20500

Statiscal discrepancy: GDP - GNI

Where GNI = GDP - net income from income

= 20600-300

= 20300

So statistical discrepancy= 20600-20300

= 300